# Actions Speak Louder than Words: Trillion-Parameter Sequential Transducers for Generative Recommendations

### v2

@@ Abstract
Large-scale recommendation systems are characterized by their reliance on high cardinality, heterogeneous features and the need to handle tens of billions of user actions on a daily basis. Despite being trained on huge volume of data with thousands of features, most Deep Learning Recommendation Models (DLRMs) in industry fail to scale with compute. Inspired by success achieved by Transformers in language and vision domains, we revisit fundamental design choices in recommendation systems. We reformulate recommendation problems as sequential transduction tasks within a generative modeling framework (“Generative Recommenders”), and propose a new architecture, HSTU, designed for high cardinality, non-stationary streaming recommendation data. HSTU outperforms baselines over synthetic and public datasets by up to 65.8% in NDCG, and is 5.3x to 15.2x faster than FlashAttention2-based Transformers on 8192 length sequences. HSTU-based Generative Recommenders, with 1.5 trillion parameters, improve metrics in online A/B tests by 12.4% and have been deployed on multiple surfaces of a large internet platform with billions of users. More importantly, the model quality of Generative Recommenders empirically scales as a power-law of training compute across three orders of magnitude, up to GPT-3/LLaMa-2 scale, which reduces carbon footprint needed for future model developments, and further paves the way for the first foundational models in recommendations.
Generative Recommenders, Machine Learning, Recommendations, Generative Modeling, Sequential Recommendations, HSTU, ICML
@@ 4.1 Validating Inductive Hypotheses of HSTU Encoder
ML-1M SASRec (2023) .2828 .5508 .7522 .1545 .2441
HSTU .3043 (+7.6%) .5728 (+4.0%) .7740 (+2.9%) .1700 (+10.1%) .2601 (+6.6%)
HSTU-large .3306 (+16.9%) .5897 (+7.1%) .7832 (+4.1%) .1858 (+20.3%) .2730 (+11.9%)
@@ 6 Conclusions
We have proposed Generative Recommenders (GRs), a new paradigm that formulates ranking and retrieval as sequential transduction tasks, allowing them to be trained in a generative manner. This is made possible by the novel HSTU encoder design, which is 5.3x-15.2x faster than state-of-the-art Transformers on 8192 length sequences, and through the use of new training and inference algorithms such as M-FALCON. With GRs, we deployed models that are 285x more complex using similar inference resources. GRs and HSTU have led to 12.4% metric improvements in production and have shown superior scaling performance compared to traditional DLRMs. Our results corroborate that user actions represent an underexplored modality in generative modeling – to echo our title, “Actions speak louder than words”.
The dramatic simplification of features in our work paves the way for the first foundational models for recommendations, search, and ads by enabling a unified feature space to be used across domains. The fully sequential setup of GRs also enables recommendation to be formulated in an end-to-end, generative setting. Both of these enable recommendation systems to better assist users holistically.
Broader Impact
We believe that our work has broad positive implications. Reducing reliance of recommendation, search, and ads systems on the large number of heterogeneous features can make these systems much more privacy friendly while improving user experiences. Enabling recommendation systems to attribute users’ long term outcomes to short-term decisions via fully sequential formulations could reduce the prevalence of content that do not serve users’ long term goals (including clickbaits and fake news) across the web, and better align incentives of platforms with user values. Finally, applications of foundational models and scaling law can help reduce carbon footprints incurred with model research and developments needed for recommendations, search, and related use cases.
@@ B.1.1 Academic Research (Traditional Sequential Recommender Settings)
The inability to capture longer term user preferences does not represent a problem due to two aspects. First, precomputed user profiles/embeddings (Xia et al., 2023) or external vector stores (Chang et al., 2023) are commonly used in modern DLRMs. Second, a significant number of contextual-, user-, and item-side features were generally employed (Zhou et al., 2018; Chen et al., 2019; Chang et al., 2023; Xia et al., 2023) and various heterogeneous networks, such as FMs (Xiao et al., 2017; Guo et al., 2017), DCNs (Wang et al., 2021), MoEs, etc. are used to transform representations and combine outputs.
In contrast to sequential settings discussed in Section B.1.1, all major industrial work defines loss over (user/request, candidate item) pairs. In the ranking setting, a multi-task binary cross-entropy loss is commonly used. In the retrieval setting, two tower setting (Covington et al., 2016) remains the dominant approach. Recent work has investigated representing the next item to recommend as a probability distribution over a sequence of (sub-)tokens, such as OTM (Zhuo et al., 2020) and DR (Gao et al., 2021) (note that in more recent work, the same setting is sometimes denoted as “generative retrieval”). They commonly utilize beam search to decode the item from sub-tokens. Advanced learned similarity functions, such as mixture-of-logits (Zhai et al., 2023a), have also been proposed and deployed as an alternative to two-tower setting and beam search given proliferation of modern accelerators such as GPUs, custom ASICs, and TPUs.
From a problem formulation perspective, we consider all work discussed above part of DLRMs (Mudigere et al., 2022) given the model architectures, features used, and losses used differ significantly from academic sequential recommender research discussed in Section B.1.1. We’d also like to point out that there have been no reports of successful applications of fully sequential ranking settings in industry, especially not at billion daily active users (DAU) scale, prior to this work.
@@ Appendix D Evaluation: Traditional Sequential Recommender Settings
ML-1M SASRec (2023) .2828 .5508 .7522 .1545 .2441
BERT4Rec .2500 (-11.6%) .1341 (-13.2%)
GRU4Rec .2811 (-0.6%) .1537 (-0.5%)
HSTU .3043 (+7.6%) .5728 (+4.0%) .7740 (+2.9%) .1700 (+10.1%) .2601 (+6.6%)
HSTU-large .3306 (+16.9%) .5897 (+7.1%) .7832 (+4.1%) .1858 (+20.3%) .2730 (+11.9%)
@@ F.2 Impact of Stochastic Length on Sequence Sparsity
Refer to captionRefer to caption
Refer to captionRefer to caption

### v3

@@ Abstract
Large-scale recommendation systems are characterized by their reliance on high cardinality, heterogeneous features and the need to handle tens of billions of user actions on a daily basis. Despite being trained on huge volume of data with thousands of features, most Deep Learning Recommendation Models (DLRMs) in industry fail to scale with compute. Inspired by success achieved by Transformers in language and vision domains, we revisit fundamental design choices in recommendation systems. We reformulate recommendation problems as sequential transduction tasks within a generative modeling framework (“Generative Recommenders”), and propose a new architecture, HSTU, designed for high cardinality, non-stationary streaming recommendation data. HSTU outperforms baselines over synthetic and public datasets by up to 65.8% in NDCG, and is 5.3x to 15.2x faster than FlashAttention2-based Transformers on 8192 length sequences. HSTU-based Generative Recommenders, with 1.5 trillion parameters, improve metrics in online A/B tests by 12.4% and have been deployed on multiple surfaces of a large internet platform with billions of users. More importantly, the model quality of Generative Recommenders empirically scales as a power-law of training compute across three orders of magnitude, up to GPT-3/LLaMa-2 scale, which reduces carbon footprint needed for future model developments, and further paves the way for the first foundation models in recommendations.
Generative Recommenders, Generative Recommendations, Sequential Recommendations, Generative Modeling, Actions Speak Louder Than Words, HSTU, Hierarchical Sequential Transduction Units, Recommendation Systems, Recommendations, Machine Learning, ICML
@@ 4.1 Validating Inductive Hypotheses of HSTU Encoder
ML-1M SASRec (2023) .2853 .5474 .7528 .1603 .2498
HSTU .3097 (+8.6%) .5754 (+5.1%) .7716 (+2.5%) .1720 (+7.3%) .2606 (+4.3%)
HSTU-large .3294 (+15.5%) .5935 (+8.4%) .7839 (+4.1%) .1893 (+18.1%) .2771 (+10.9%)
@@ 6 Conclusions
We have proposed Generative Recommenders (GRs), a new paradigm that formulates ranking and retrieval as sequential transduction tasks, allowing them to be trained in a generative manner. This is made possible by the novel HSTU encoder design, which is 5.3x-15.2x faster than state-of-the-art Transformers on 8192 length sequences, and through the use of new training and inference algorithms such as M-FALCON. With GRs, we deployed models that are 285x more complex while using less inference compute. GRs and HSTU have led to 12.4% metric improvements in production and have shown superior scaling performance compared to traditional DLRMs. Our results corroborate that user actions represent an underexplored modality in generative modeling – to echo our title, “Actions speak louder than words”.
The dramatic simplification of features in our work paves the way for the first foundation models for recommendations, search, and ads by enabling a unified feature space to be used across domains. The fully sequential setup of GRs also enables recommendation to be formulated in an end-to-end, generative setting. Both of these enable recommendation systems to better assist users holistically.
Impact Statements
We believe that our work has broad positive implications. Reducing reliance of recommendation, search, and ads systems on the large number of heterogeneous features can make these systems much more privacy friendly while improving user experiences. Enabling recommendation systems to attribute users’ long term outcomes to short-term decisions via fully sequential formulations could reduce the prevalence of content that do not serve users’ long term goals (including clickbaits and fake news) across the web, and better align incentives of platforms with user values. Finally, applications of foundation models and scaling law can help reduce carbon footprints incurred with model research and developments needed for recommendations, search, and related use cases.
@@ B.1.1 Academic Research (Traditional Sequential Recommender Settings)
Despite using short sequence lengths, most DLRMs can successfully capture long-term user preferences. This can be attributed to two key aspects. First, precomputed user profiles/embeddings (Xia et al., 2023) or external vector stores (Chang et al., 2023) are commonly used in modern DLRMs, both of which effectively extend lookback windows. Second, a significant number of contextual-, user-, and item-side features were generally employed (Zhou et al., 2018; Chen et al., 2019; Chang et al., 2023; Xia et al., 2023) and various heterogeneous networks, such as FMs (Xiao et al., 2017; Guo et al., 2017), DCNs (Wang et al., 2021), MoEs, etc. are used to transform representations and combine outputs.
In contrast to sequential settings discussed in Section B.1.1, all major industrial work defines loss over (user/request, candidate item) pairs. In the ranking setting, a multi-task binary cross-entropy loss is commonly used. In the retrieval setting, two tower setting (Covington et al., 2016) remains the dominant approach. Recent work has investigated representing the next item to recommend as a probability distribution over a sequence of (sub-)tokens, such as OTM (Zhuo et al., 2020), and DR (Gao et al., 2021) (note that in other recent work, the same setting is sometimes denoted as “generative retrieval”). They commonly utilize beam search to decode the item from sub-tokens. Advanced learned similarity functions, such as mixture-of-logits (Zhai et al., 2023a), have also been proposed and deployed as an alternative to two-tower setting and beam search given proliferation of modern accelerators such as GPUs, custom ASICs, and TPUs.
From a problem formulation perspective, we consider all work discussed above part of DLRMs (Mudigere et al., 2022) given the model architectures, features used, and losses used differ significantly from academic sequential recommender research discussed in Section B.1.1. It’s also worth remarking that there have been no successful applications of fully sequential ranking settings in industry, especially not at billion daily active users (DAU) scale, prior to this work.
@@ Appendix D Evaluation: Traditional Sequential Recommender Settings
ML-1M SASRec (2023) .2853 .5474 .7528 .1603 .2498
BERT4Rec .2843 (-0.4%) .1537 (-4.1%)
GRU4Rec .2811 (-1.5%) .1648 (+2.8%)
HSTU .3097 (+8.6%) .5754 (+5.1%) .7716 (+2.5%) .1720 (+7.3%) .2606 (+4.3%)
HSTU-large .3294 (+15.5%) .5935 (+8.4%) .7839 (+4.1%) .1893 (+18.1%) .2771 (+10.9%)